# Consecutive Model Editing with Batch alongside HooK Layers

### v1

@@ Abstract
As the typical retraining paradigm is unacceptably time- and resource-consuming, researchers are turning to model editing in order to seek an effective, consecutive, and batch-supportive way to edit the model behavior directly. Despite all these practical expectations, existing model editing methods fail to realize all of them. Furthermore, the memory demands for such succession-supportive model editing approaches tend to be prohibitive, frequently necessitating an external memory that grows incrementally over time. To cope with these challenges, we propose COMEBA-HK, a model editing method that is both consecutive and batch-supportive. COMEBA-HK is memory-friendly as it only needs a small amount of it to store several hook layers with updated weights. Experimental results demonstrate the superiority of our method over other batch-supportive model editing methods under both single-round and consecutive batch editing scenarios. Extensive analyses of COMEBA-HK have been conducted to verify the stability of our method over 1) the number of consecutive steps and 2) the number of editing instances. Our code will be released via https://github.com/anonymous.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) (Chung et al., 2022; OpenAI, 2023; Black et al., 2022; Touvron et al., 2023) have been demonstrated to be capable of recalling factual knowledge about the real world (Brown et al., 2020; Petroni et al., 2020). Nevertheless, researchers also reveal that LLMs often fail to recall the most up-to-date knowledge or information and some specialized knowledge if they are not periodically updated (Liska et al., 2022; Agarwal & Nenkova, 2022; Lazaridou et al., 2021). Despite the fact that fresh and customizable knowledge is highly desired in many areas, such as text generation, question-answering, reasoning, etc., updating the model via retraining is both time- and resource-consuming. Additionally, researchers have uncovered that well-trained LLMs do make mistakes. One popular sort of mistake is called hallucination (Tonmoy et al., 2024), which means that LLMs generate text based on ”hallucinated” fake knowledge. Although many researchers have tried to mitigate this issue (Qiu et al., 2023; Mündler et al., 2023; Kang et al., 2023; Varshney et al., 2023), the strategy to fix this bug remains unclear. Therefore, researchers have started to seek an efficient approach that could edit LLMs in a customizable, cost-effective way.
@@
To this end, recent years have witnessed many efforts in investigating the model-editing techniques to bypass the retraining paradigm and edit the LLMs directly (Meng et al., 2022; Hartvigsen et al., 2022; Li et al., 2023; Mitchell et al., 2022a, b). Accordingly, several new datasets (e.g., ZsRE (Levy et al., 2017) and COUNTERFACT (Meng et al., 2023)) and corresponding metrics (e.g., reliability, generality, locality, portability (Yao et al., 2023)) are proposed to facilitate the development in this field. However, these methods either require extra expensive training of a meta-network (Mitchell et al., 2022a), or a classifier (Mitchell et al., 2022b), which causes time and resources overhead, or demands an external memory of explicit edit instances for reference (Mitchell et al., 2022b), which inevitably escalates the memory requirement. Further, most of the methods proposed so far are only targeted for single-round editing, where the model is rolled back to the initial state after each edit step. We believe this deviates from the application scenario in reality since most users anticipate an editing approach that is succession-supportive and batch-supportive, i.e., a method that advocates consecutive batch editing.
@@
In light of these issues, we propose a novel method, named COMEBA-HK, which performs COnsecutive Model Editing with BAtch alongside HooK layers. Specifically, COMEBA-HK supports consecutive batch editing and utilizes the hook layers to separate weight change from the original model weight. COMEBA-HK does not need any training or large explicit external memory that stores editing instances. It only needs a tiny amount of memory to collect the optimized weight in the hook layer. To achieve this, we propose a new transformer memory updating mechanism that supports consecutive batch editing settings and design a simple yet effective local editing scope identification technique used in the hook layer that can accurately detect the inputs in the local editing scope. We demonstrate the effectiveness of our method via extensive experiments on ZsRE and COUNTERFACT datasets using two popular autoregressive language models, GPT2-XL and GPT-J (6B). Both the single-round batch settings and consecutive batch settings are included, with the number of consecutive batch editing instances ranging from 1k to 10k. An analysis of the editing scope identification has also been conducted to validate the method. Beyond all these, we implement comprehensive ablation studies to verify the validity of each component and discuss the optimal parameter settings in the method.
@@ 2 Preliminaries of Model Editing
$f_{\theta^{\prime}}(x)=\left\{\begin{array}[]{l}y_{t}\quad\quad~{}~{}\text{if}~{}x\in\mathcal{X}_{t}\\ f_{\theta}(x)\quad\text{if}~{}x\notin\mathcal{X}_{t}\end{array}\right..$ (1)
In particular, there are three standard criteria for model editing, namely Reliability, Generality, and Locality (Yao et al., 2023; Mitchell et al., 2022a, b). Suppose the prediction of the original model to the statement “What is the native language of Joe Biden?” is “French”, and the expected post-edit model prediction is “English”. To verify the Reliability, we use the same original statement as input and then assess whether the post-edit model predicts “English” as desired. For Generality, a rephrased statement “What is the native language of Joe Biden?” could be inputted into the edited model to assess whether the output of the model remains as “English”. Locality suggests that the model output of an irrelevant statement like “What is the native language of Donald Trump?” should remain unaffected, which means that the post-edit model should output whatever the initial model output to this statement.
The current problem settings of model editing can be generally categorized into three groups (Yao et al., 2023):
@@
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ .$ (2)
@@
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}\sum\nolimits_{t=1}^{n}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ .$ (3)
where $n\leq\mid{\mathcal{X}_{t}}\mid$ is the batch size and it varies for different methods (Meng et al., 2023; Mitchell et al., 2022a, b; Meng et al., 2022).
3) Sequential Editing requires every single edit to be performed successively, and evaluation has to be conducted after a series of knowledge updates (Hartvigsen et al., 2022):
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}\sum\nolimits_{t=1}^{\mid{\mathcal{X}_{t}}\mid}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ .$ (4)
In this work, we investigate a new and more practical problem setting for model editing, namely Consecutive Batch Editing, which aims at executing the editing in a consecutive and batch-edits-supportive way.
$\theta^{\prime}\leftarrow\operatorname*{argmin}_{\theta}\sum_{s=0}^{\lceil\mid{\mathcal{X}_{t}}\mid/n\rceil}\sum_{t=s\times n}^{\min((s+1)\times n,\mid{\mathcal{X}_{t}}\mid)}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)\ ,$ (5)
@@ 3 Method
We first discuss our method under the single-layer consecutive batch editing setting. Explicitly, we first discuss the process of extending the single-layer updating mechanism in MEMIT (Meng et al., 2023) from a scenario of single-round batch editing to consecutive batch editing. Then, we describe the motivation, definition, and application of the hook layer and the local editing scope identification operation employed in the hook layer. The practicality of the operation is also clarified. Finally, the motivation and procedure for broadening the method from the single-layer case to the multi-layer case are discussed.
@@ 3.1.1 Batch Editing Menchanism
Meng et al. (2023) demonstrate an effective single-layer editing method using minimal squared error. Although it supports multiple edits on a single round, the updates do not account for scenarios involving consecutive updates. In this section, we extend this approach to include consecutive scenarios. Strang (2022) suggests Eq.6 can be optimized if we can find a $W_{0}$ that satisfies Eq.7, where $K_{0}=[k_{1}|k_{2}|...|k_{n}]$ and $V_{0}=[v_{1}|v_{2}|...|v_{n}]$ are set of key-value pairs we want to store into the layer.
$W_{0}=\operatorname*{argmin}_{W}\sum\nolimits_{i=1}^{n}\norm{Wk_{i}-v_{i}}^{2}$ (6)
$W_{0}K_{0}K_{0}^{T}=V_{0}K_{0}^{T}\ .$ (7)
Thanks to the well-conducted pre-training procedure for most of the available transformer-based LLMs, we can assume that the pre-training weight $W_{0}$ satisfies Eq.7, i.e., serves as the optimal solution for Eq.6.
@@
$W_{1}=\operatorname*{argmin}_{W}(\sum_{i=1}^{r}\norm{Wk_{i}-v_{i}}^{2}+\sum_{i=r+1}^{r+u}\norm{Wk_{i}-v_{i}}^{2})$ (8)
$W_{1}[K_{1}\,K_{2}][K_{1}\,K_{2}\,]^{T}\,=\,[V_{1}\,V_{2}][K_{1}\,K_{2}]^{T}\ ,$ (9)
where $K_{1}=[k_{1}|k_{2}|\dots|k_{r}](r\geq n)$ and $V_{1}=[v_{1}|v_{2}|\dots|v_{r}]$ is the set of key-value pairs that have been updated and $K_{2}=[k_{r+1}|k_{r+2}|...|k_{r+u}]$ and $V_{2}=[v_{r+1}|v_{r+2}|...|v_{r+u}]$ is the set of key-values that are going to be edited. Therefore, the objective (Eq.8) indicates that we want an optimal $W_{1}$ that successfully updates the new associations while maintaining the old key-value pairs.
@@
In a real consecutive editing scenario, $r$ increases and starts with $n$, and each batch-editing iteration is optimized through the objective (Eq.8). Hence, we can conclude that $W_{1}K_{1}K_{1}^{T}=V_{1}K_{1}^{T}$. Subtracting it from Eq.11, we get:
@@
$\Delta=RK_{2}^{T}C_{accu}^{-1}\ ,$ (13)
where $R=(V_{2}-W_{1}K_{2})$ is the residual error evaluated on the most recent updated weights (Note that $W_{1}=W_{0}$ if and only if $r=n$) and $C_{accu}=(K_{1}K_{1}^{T}+K_{2}K_{2}^{T})$ the accumulation sum of editing keys’ outer product. Note that $K_{1}K_{1}^{T}$ is computed via:
$K_{1}K_{1}^{T}=K_{0}K_{0}^{T}+K^{\prime}K^{\prime T}\ ,$ (14)
where $K_{0}$ is the set of pre-training keys that have been contained in the pre-training weight, $K^{\prime}=[k_{n+1}|k_{n+2}|...|k_{r}]$ denotes the updated keys proceeding to current editing step. We follow (Meng et al., 2023) to model $K_{0}K_{0}^{T}$ as the uncentered covariance of some randomly sampled inputs:
$K_{0}K_{0}^{T}=\lambda E[kk^{T}]\ .$ (15)
Note that the $\lambda$ represents a factor that balances the pre-trained and the whole updated associations. We follow the definitions and applications of keys and values in (Meng et al., 2023, 2022), where keys are the activations at the last token of the subject and values are gradient-descent optimized vectors that maximize the model’s prediction for the target object (Fig.1).
Figure 1: Single layer update with hook layer (residual connections are omitted).
@@ 3.1.2 Hook Layer
Yao et al. (2023) demonstrated that those editing methods that directly modify the model parameter in place struggle with sequential editing. Specifically, the locality decreases drastically when the number of iterations increases. Meanwhile, those methods that freeze the model parameters show more stable performance over iterations. This indicates that it might be helpful to separate the editing change from the model itself. However, it is too memory-costly to straightforwardly make another copy of the model or apply an external memory (Mitchell et al., 2022b; Hartvigsen et al., 2022) for a consecutive batch editing scenario. Therefore, a potentially cost-effective trade-off might be only copying part of the model layers and editing merely on those copied layer weights. This allows us to have access to both the original and edited outputs so that we can design criteria to precisely swap specific parts of the original output for the edited outputs whenever necessary.
With all these motivations, we introduce the hook layer (Fig.1), which takes the original model layer weights as the weight initialization and is responsible for all editing weight alteration in the whole editing process of COMEBA-HK. It serves a similar idea with the forward hook function defined in popular ML libraries like PyTorch, which adjusts the original forward layer output based on some predefined criterion. Theoretically, the hook layer can be hung up on any target linear layer in the transformer, Nevertheless, we mainly focus on the critical path identified in (Meng et al., 2023, 2022) as they are verified to be crucial for fact associations storage in the autoregressive language model.
As shown in Fig.1, there are generally two sorts of hook layers in this work, namely, the Temporary hook layer and the Validated hook layer. The temporary hook layer is mainly applied during the process of editing. It replaces the whole original output with the output from the hook layer so that the residual with respect to the hook layer weights instead of the original layer weights is computed. The hook layer weights are then updated (Eq.13) using the calculated residual and the accumulated sum of keys’ outer product. Validated hook layers are employed after each batch editing step at the layer, and they inherit the updated weights from the temporary hook layer. We describe the operation conducted by it in the section 3.1.3.
Figure 2: Multiple layer update with hook layer (Attention module and the first FFN are omitted).

### v3

# First Author Affiliation / Address line 1 Affiliation / Address line 2

@@ Abstract
As the typical retraining paradigm is unacceptably time- and resource-consuming, researchers are turning to model editing to find an effective way that supports both consecutive and batch scenarios to edit the model behavior directly. Despite all these practical expectations, existing model editing methods fail to realize all of them. Furthermore, the memory demands for such sequential model editing approaches tend to be prohibitive, frequently necessitating an external memory that grows incrementally over time. To cope with these challenges, we propose CoachHooK, a model editing method that simultaneously supports sequential and batch editing. CoachHooK is memory-friendly as it only needs a small amount of it to store several hook layers whose size remains unchanged over time. Experimental results demonstrate the superiority of our method over other batch-supportive model editing methods under both single-round and consecutive batch editing scenarios. Extensive analyses of CoachHooK have been conducted to verify the stability of our method over a number of consecutive steps. Corresponding author.
Consecutive Batch Model Editing with HooK Layers thanks:   This research was substantially supported by the Center for Perceptual and Interactive Intelligence (CPII) Ltd. under the Innovation and Technology Commission’s InnoHK scheme.
Shuaiyi Li1, Yang Deng2,†, Deng Cai3, Hongyuan Lu1, Liang Chen1, Wai Lam1 1 The Chinese University of Hong Kong, 2 Singapore Management University, 3 Tencent AI Lab sli@se.cuhk.edu.hk, ydeng@smu.edu.sg, thisisjcykcd@gmail.com, 1155167604@link.cuhk.edu.hk, lchen@se.cuhk.edu.hk, wlam@se.cuhk.edu.hk
@@ 1 Introduction
Large Language Models (LLMs) Chung et al. (2022); OpenAI (2023); Black et al. (2022); Touvron et al. (2023) have been demonstrated to be capable of recalling factual knowledge about the real world Brown et al. (2020); Petroni et al. (2020). Nevertheless, researchers also reveal that LLMs often fail to recall the most up-to-date knowledge or information and some specialized knowledge if they are not periodically updated Liska et al. (2022); Agarwal and Nenkova (2022); Lazaridou et al. (2021). Despite the fact that fresh and customizable knowledge is highly desired in many areas, such as text generation, question-answering, reasoning, etc., updating the model via retraining is both time- and resource-consuming. Additionally, researchers have uncovered that well-trained LLMs do make mistakes. One popular sort of mistake is called hallucination Tonmoy et al. (2024), which means that LLMs generate text based on "hallucinated" fake knowledge. Although many researchers have tried to mitigate this issue Qiu et al. (2023); Mündler et al. (2023); Kang et al. (2023); Varshney et al. (2023), the strategy to fix this bug remains unclear. Therefore, researchers have started to seek an efficient approach that could edit LLMs in a customizable, cost-effective way.
@@
To this end, recent years have witnessed many efforts in investigating the model-editing techniques to bypass the retraining paradigm and edit the LLMs directly Meng et al. (2022); Hartvigsen et al. (2022); Li et al. (2023b); Mitchell et al. (2022a, b). Accordingly, several new datasets (e.g., ZsRE Levy et al. (2017) and COUNTERFACT Meng et al. (2023)) and corresponding metrics (e.g., reliability, generality, locality, portability Yao et al. (2023)) are proposed to facilitate the development in this field. However, these methods either require extra expensive training of a meta-network Mitchell et al. (2022a); De Cao et al. (2021), or a classifier Mitchell et al. (2022b), which causes time and resources overhead, or demands an external memory of explicit edit instances for reference Mitchell et al. (2022b); Hartvigsen et al. (2022), which inevitably escalates the memory requirement. Further, most of existing methods are only evaluated on single-round editing, where the model is rolled back to the initial state after each edit step. This deviates from the application scenario in reality since most users anticipate an editing approach that allows sequential and batch editing.
@@
In light of these issues, we propose a novel method, named CoachHooK [^1] Code is released at https://github.com/Syon-Li/CoachHook., which performs Consecutive Batch Model Editing with HooK layers. Specifically, CoachHooK supports consecutive batch editing and utilizes the hook layers to separate weight change from the original model weight. CoachHook does not need any training on parameters or large explicit external memory that stores editing instances. It only needs a reasonable amount of memory to collect the optimized weight in the hook layer. To achieve this, we propose a new transformer memory updating mechanism that supports consecutive batch editing settings and design a simple yet effective local editing scope identification technique used in the hook layer that can accurately detect the inputs in the local editing scope. We demonstrate the effectiveness of our method via extensive experiments on ZsRE and COUNTERFACT datasets using two popular autoregressive language models, GPT2-XL and GPT-J (6B). Both the single-round batch settings and consecutive batch settings are included, with the total number of editing instances ranging from 1k to 10k. An analysis of the editing scope identification has also been conducted to validate the method. Beyond all these, we implement comprehensive ablation studies to verify the validity of each component and discuss the optimal hyperparameter settings in the method.
@@ 2 Preliminaries of Model Editing
$f_{\theta^{\prime}}(x)=\left\{\begin{array}[]{l}y_{t}\quad\quad~{}~{}\text{if}~{}x\in\mathcal{X}_{t}\\ f_{\theta}(x)\quad\text{if}~{}x\notin\mathcal{X}_{t}\end{array}\right.$ (1)
In particular, there are three standard criteria for model editing, namely Reliability, Generality, and Locality Yao et al. (2023); Mitchell et al. (2022a, b). Suppose the prediction of the original model to the prompt "What is the native language of Joe Biden?" is "French", and the expected post-edit model prediction is "English". To verify the Reliability, we use the same original prompt as input and then assess whether the post-edit model predicts "English" as desired. For Generality, a rephrased prompt "The mother tongue of Joe Biden is" could be inputted into the edited model to assess whether the output of the model remains as "English". Locality suggests that the model output of an irrelevant prompt like "What is the native language of Donald Trump?" should remain unaffected, which means that the post-edit model should output whatever the initial model output to this prompt.
The current problem settings of model editing can be generally categorized into three groups Yao et al. (2023):
@@
$\theta^{\prime}\leftarrow\operatorname*{argmin}\nolimits_{\theta}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)$ (2)
@@
$\theta^{\prime}\leftarrow\operatorname*{argmin}\nolimits_{\theta}\sum\nolimits_{t=1}^{n}(\parallel{f_{\theta}(x_{t})-y_{t}}\parallel)$ (3)
where $n\leq\mid{\mathcal{X}_{t}}\mid$ is the batch size and it varies for different methods Meng et al. (2023); Mitchell et al. (2022a, b); Meng et al. (2022).
3) Sequential Editing requires every single edit to be performed successively, and evaluation has to be conducted after a series of knowledge updates Hartvigsen et al. (2022):
${\theta_{\mid{\mathcal{X}_{t}}\mid}}^{\prime}\leftarrow\operatorname*{argmin}\nolimits_{\theta_{t}}\sum\nolimits_{t=1}^{\mid{\mathcal{X}_{t}}\mid}(\parallel{f_{\theta_{t}}(x_{t})-y_{t}}\parallel)$ (4)
In this work, we investigate a new and more practical setting for model editing, namely Consecutive Batch Editing, which aims at executing the editing in a consecutive batch editing way:
$\begin{split}{\theta_{\lceil\mid{\mathcal{X}_{t}}\mid/n\rceil}}^{\prime}&\leftarrow\operatorname*{argmin}_{\theta_{s}}\sum_{s=0}^{\lceil\mid{\mathcal{X}_{t}}\mid/n\rceil}\Big{[}\\ &\sum_{t=s\cdot n}^{\min((s+1)\cdot n,\mid{\mathcal{X}_{t}}\mid)}(\parallel{f_{\theta_{s}}(x_{t})-y_{t}}\parallel)\Big{]}\end{split}$ (5)
@@ 3 Method
We first discuss our method under the single-layer consecutive batch editing setting. Explicitly, we first discuss the process of extending the single-layer updating mechanism in MEMIT Meng et al. (2023) from a scenario of single-round batch editing to consecutive batch editing. Then, we introduce the hook layer and the local editing scope identification operation employed in the hook layer. The practicality of the operation is also clarified. Finally, we broaden the method from single-layer to multi-layer scenarios.
@@ 3.1.1 Batch Editing Menchanism
Meng et al. (2023) demonstrate an effective single-layer editing method using minimal squared error. Although it supports multiple edits on a single round, the updates do not account for scenarios involving consecutive updates. In this section, we extend this approach to include consecutive scenarios. Following Meng et al. (2023, 2022), we analyse the model layer weights $W_{0}$ as a linear associative memory Kohonen (1972a); Anderson (1972a) that stores associations between a set of keys $k_{i}$ and values $v_{i}$ using minimal squared error:
$W_{0}=\operatorname*{argmin}\nolimits_{W}\sum\nolimits_{i=1}^{n}\norm{Wk_{i}-v_{i}}^{2}$ (6)
In this work, $W_{0}$ is the weight of the second layer of the model’s FFN part (denoted as $W_{proj}^{l}$). For simplicity, we stack keys and values into matrices $K_{0}=[k_{1}|k_{2}|...|k_{n}]$ and $V_{0}=[v_{1}|v_{2}|...|v_{n}]$, then Eq. 6 can be optimized by solving Strang (2022):
$W_{0}K_{0}K_{0}^{T}=V_{0}K_{0}^{T}$ (7)
Thanks to the well-conducted pre-training procedure for most of the available LLMs, we can assume that the pre-trained weight $W_{0}$ satisfies Eq. 7, i.e., serves as the optimal solution for Eq. 6.
@@
$\begin{split}\hat{W_{1}}=\operatorname*{argmin}\nolimits_{W}(&\sum\nolimits_{i=1}^{r}\norm{Wk_{i}-v_{i}}^{2}\\ &+\sum\nolimits_{i=r+1}^{r+u}\norm{Wk_{i}-v_{i}}^{2})\end{split}$ (8)
Following Eq. 7, we conclude that Eq. 8 can be optimized if we can solve:
$\hat{W_{1}}[K_{1}\,K_{2}][K_{1}\,K_{2}\,]^{T}\,=\,[V_{1}\,V_{2}][K_{1}\,K_{2}]^{T}$ (9)
where $K_{1}=[k_{1}|k_{2}|\dots|k_{r}](r\geq n)$ and $V_{1}=[v_{1}|v_{2}|\dots|v_{r}]$ is the set of key-value pairs that have been updated and $K_{2}=[k_{r+1}|k_{r+2}|...|k_{r+u}]$ and $V_{2}=[v_{r+1}|v_{r+2}|...|v_{r+u}]$ is the set of key-values that are going to be edited. Therefore, the objective (Eq. 8) indicates that we want an optimal $\hat{W_{1}}$ that successfully updates the new associations while maintaining the old key-value pairs.
@@
The $\Delta$ means the desired weight change to update the new associations $K_{2},V_{2}$ and $W_{1}$ is the weight that has been updated for the associations $K_{1},V_{1}$ (Note that $W_{1}=W_{0}$ if and only if $r=n$). In a real consecutive editing scenario, $r$ increases and starts with $n$, and each batch-editing iteration is optimized through the objective (Eq. 8). Hence, we can conclude that $W_{1}K_{1}K_{1}^{T}=V_{1}K_{1}^{T}$. Subtracting it from Eq. 11, we get:
@@
Further rearranging it, we have:
$\Delta=RK_{2}^{T}C_{accu}^{-1}$ (13)
where $R=(V_{2}-W_{1}K_{2})$ is the residual error evaluated on the most recent updated weights. $C_{accu}=(K_{1}K_{1}^{T}+K_{2}K_{2}^{T})$ is the accumulation sum of editing keys’ outer product, and we have
$K_{1}K_{1}^{T}=K_{0}K_{0}^{T}+K^{\prime}K^{\prime T}$ (14)
where $K_{0}$ is the set of pre-training keys that have been contained in the pre-training weight, $K^{\prime}=[k_{n+1}|k_{n+2}|...|k_{r}]$ denotes the updated keys proceeding to current editing step. We follow Meng et al. (2023) to model $K_{0}K_{0}^{T}$ as the uncentered covariance of some randomly sampled inputs:
$K_{0}K_{0}^{T}=\lambda E[kk^{T}]$ (15)
Note that the $\lambda$ represents a factor that balances the pre-trained and the whole updated associations. We follow the definitions of keys and values in Meng et al. (2023, 2022), where keys are the activations at the last token of the subject (such as "Joe Biden" for example provided in §1) and values are gradient-descent optimized vectors that maximize the model’s prediction for the target object.
Figure 1: Single layer update with hook layer (residual connections are omitted). $\parallel$.$\parallel$ . ∥ means calculate the L2-norm over the keys’ dimension ($m$). For each updating of a single batch edits, the temporary hook layer is used at the beginning to ensure $\Delta$ is computed based on $W_{h}^{l}$. After the weights update, the validated hook layer is applied to determine whether to use the original layer or hook layer for each token. This process can be implemented iteratively to support consecutive batch editing. Note that the temporary hook layer weight of a new iteration is copied from the validated hook layer weight of the previous iteration. So, the validated hook layer keeps track of the updated layer from previous edits by retaining the weight from the previous iteration.
@@ 3.1.2 Hook Layer
Yao et al. (2023) demonstrate that those editing methods that directly modify the model parameter in place struggle with sequential editing. Specifically, the locality decreases drastically when the number of iterations increases. Meanwhile, those methods that freeze the model parameters show more stable performance over iterations. This indicates that it might be helpful to separate the editing change from the model itself. However, directly applying an external memory Mitchell et al. (2022b); Hartvigsen et al. (2022) that grows over time for a consecutive batch editing scenario is too memory-costly. Therefore, we aim to seek an approach that could store associations without regularly increasing external memory while preserving the original model parameters.
In light of these motivations, we introduce the hook layer (Fig. 1), which takes the original model layer weights as the weight initialization and is responsible for all editing weight alteration in the whole editing process of CoachHooK. It is similar to the forward hook function defined in popular ML libraries like PyTorch, which adjusts the original forward layer output based on predefined criteria. Theoretically, the hook layer can be hung on any target linear layer in the transformer. Nevertheless, we mainly focus on the critical path identified in Meng et al. (2023, 2022) as they are verified to be crucial for fact association storage in the autoregressive language model.
As shown in Fig. 1, there are generally two sorts of hook layers in this work, namely, the Temporary hook layer and the Validated hook layer. The temporary hook layer is temporarily applied during the weight-updating process. It replaces the original output with the output from the hook layer so that the residual is computed based on the hook layer weight. The hook layer weights are then updated (Eq. 13) using the calculated residual and the accumulated sum of the keys’ outer product. Validated hook layers inherit the updated weights from the temporary hook layer, and are employed after each single-batch weight updating process at the layer.